ou have just made your first $5,400 contribution to your retirement account. Assume you earn a return of 13 percent per year and make no additional contributions.

  

a.

What will your account be worth when you retire in 41 years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

b. What if you wait 10 years before contributing? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
a. Future value for starting now
b. Future value if waiting 10 years
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1.A=$5400*(1.13)^41

=$5400*150.0431533

=$810,233.03(Approx).

2.A=$5400*(1.13)^31

=$5400*44.20096469

=$238,685.21(Approx).
Answer #2

SOLUTION :


a. Future value starting now =  $810,233.03 (ANSWER)


Work :


FV available on retirement after 41 years :

= Amount invested ( 1 + r)^n 

= 5400 (1 + 0.13)^41

= 810233.03 ($)  




b.  Future value if waiting 10 years = $238,685.21 (ANSWER)


Work :


If deposit of $5,500 is made after 10 years from now,

FV available at retirement after 41 years :

= Amount invested ( 1 + r)^n 

= 5400 (1 + 0.13)^(41 - 10)

= 238685.21 ($)  

answered by: Tulsiram Garg